ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 244, 2021

Algom-Kfir, Y., A. Hilion and E. Stark The visual boundary of hyperbolic free-by-cyclic groups

501

Amarra, C., L. Morgan and C. E. Praeger Generalised shuffle groups

807

Arias-de-Reyna, S., F. Legrand and G. Wiese Galois families of modular forms and application to weight one

1

Aroca, J. The conjugacy problem for symmetric Thompson-like groups

49

Baudier, F. P., G. Lancien, P. Motakis and Th. Schlumprecht The geometry of Hamming-type metrics and their embeddings into Banach spaces

681

Berkowitz, R., A. Sah and M. Sawhney Number of arithmetic progressions in dense random subsets of ℤ/nℤ

589

Brough, J., Y. Liu and A. Paolini The block graph of a finite group

293

Buss, A. and A. Sims Opposite algebras of groupoid C^{∗}-algebras  759

Caspers, M., F. Sukochev and D. Zanin Weak (1, 1) estimates for multiple operator integrals and generalized absolute value functions

245

Dolfi, S., E. Pacifici, L. Sanus and V. Sotomayor Groups whose prime graph on class sizes has a cut vertex

775

Feldheim, N. D. and O. N. Feldheim Mean and minimum of independent random variables

857

Feldheim, O. N. See N. D. Feldheim Galkowski, J. and S. Zelditch Lower bounds for Cauchy data on

curves in a negatively curved surface

971

Grafakos, L. An improvement of the Marcinkiewicz multiplier theorem

163

Greenberg, N., J. S. Miller and A. Nies Highness properties close to PA completeness

419

Gruber, D. and J. M. Mackay Random triangular Burnside groups  75

Hilion, A. See Y. Algom-Kfir Hoang, H. P. and T. Mütze Combinatorial generation via permu-

tation languages. II. Lattice congruences

359

Hu, H. and N. Hu Double-bosonization and Majid’s Conjecture (II): irregular R-matrices and type-crossing to G_{2} 901 Hu, N. See H. Hu Jahel, C. and A. Zucker Minimal model-universal flows for locally

compact Polish groups

743

Kazaniecki, K., P. Pakosz and M. Wojciechowski An invariant version of the little Grothendieck theorem for Sobolev spaces

33

Keller, C. and S. Smorodinsky A new lower bound on Hadwiger– Debrunner numbers in the plane

649

Kifer, Y. and F. Yang Geometric law for numbers of returns until a hazard under ϕ-mixing

319

Lancien, G. See F. P. Baudier Legrand, F. See S. Arias-de-Reyna Liu, Y. See J. Brough

Lofano, D. and A. Newman The worst way to collapse a simplex  625

Mackay, J. M. See D. Gruber

Malcolm, A. J. On the p-width of finite simple groups  127

Martino, A., M. C. H. Tointon, M. Valiunas and E. Ventura Probabilistic nilpotence in infinite groups

539

Miller, J. S. See N. Greenberg Morgan, L. See C. Amarra Motakis, P. See F. P. Baudier Mütze, T. See H. P. Hoang

Navas, A. (Un)distorted diffeomorphisms in different regularities  727

Newman, A. See D. Lofano Nies, A. See N. Greenberg Olukoya, F. Automorphism towers of groups of homeomorphisms of

Cantor space

883

Pacifici, E. See S. Dolfi Pakosz, P. See K. Kazaniecki

Palacín, D. Finite groups contain large centralizers  621

Paolini, A. See J. Brough Pascoe, J. E. The outer spectral radius and dynamics of completely

positive maps

945

Passman, D. S. and L. W. Small Finite presentation  185

Praeger, C. E. See C. Amarra Sah, A. See R. Berkowitz Sanus, L. See S. Dolfi Sasane, A. On the existence of spatially tempered null solutions to

linear constant coefficient PDEs

273

Sawhney, M. See R. Berkowitz Schlumprecht, Th. See F. P. Baudier Sims, A. See A. Buss Small, L. W. See D. S. Passman Smorodinsky, S. See C. Keller Sotomayor, V. See S. Dolfi Stark, E. See Y. Algom-Kfir Sukochev, F. See M. Caspers Tang, X., J. J. Zhang and X.-G. Zhao Cancellation of Morita and

skew types

467

Thakur, M. On R-triviality of F_{4}  145

Tointon, M. C. H. See A. Martino Valiunas, M. See A. Martino Ventura, E. See A. Martino Wiese, G. See S. Arias-de-Reyna Wojciechowski, M. See K. Kazaniecki Yang, F. See Y. Kifer

Yavicoli, A. Patterns in thick compact sets  95

Yom Din, A. Second adjointness for tempered admissible representations of a real group

215

Zanin, D. See M. Caspers Zelditch, S. See J. Galkowski Zhang, J. J. See X. Tang Zhao, X.-G. See X. Tang Zucker, A. See C. Jahel